# Quantum adiabatic optimization with Rydberg arrays: localization phenomena and encoding strategies

### v1

@@ III.2 Crossing-with-edge setup
Figure 4: Scaling of the minimum gap $\delta E_{m}$ along the standard protocol for the crossing-with-edge setup. $\delta E_{m}$ (normalized to the maximum Rabi frequency $\Omega_{0}$) as a function of the number of atoms per vertex-wire $N$, for the configurations: (a) $m_{x}=m_{y}=N/2$, (b) $m_{x}=N/2$, $m_{y}=4$, (c) $m_{x}=m_{y}=4$, and (d) $m_{x}=m_{y}=N-4$. For symmetric geometries (a,c,d), the ground state at final time is $|01\rangle_{l}$. In the asymmetric geometry (b), $|10\rangle_{l}$ (blue) and $|01\rangle_{l}$ (orange) are considered separately. In the top-right inset of (c), the gap $\delta E$ is shown as a function of $\Delta/\Omega$, for $N$ ranging from $8$ (light blue) to $48$ (dark blue).
Two vertex-wires intersecting via the CWE gadget [see Fig. 1(d)] encode the simple graph consisting of two vertices connected by an edge. Even though the two vertices play a symmetric role in the original graph, this is not necessarily the case for the two vertex-wires if $m_{x}\neq m_{y}$. We thus consider, for a given geometry (i.e., for a given choice of $m_{x}$ and $m_{y}$), two distinct MWIS problems, where $\delta_{x}>\delta_{y}$ and $\delta_{x}<\delta_{y}$. The solution to these problems is encoded in the $|10\rangle_{l}$ and $|01\rangle_{l}$ states of the two vertex-wires, respectively. Motivated by the possible positions of the gadget in a large network, we analyze four different geometrical configurations with $N_{x}=N_{y}=N$. We consider the gadget placed: (a) at the midpoint of the two vertex-wires (symmetric case, ${m_{x}=m_{y}=N/2}$); (b) close to the beginning of the $y$ vertex-wire ($m_{x}=N/2$, $m_{y}=c=4$); (c) close to the beginning of both vertex-wires ($m_{x}=m_{y}=c=4$); (d) close to the end of both vertex-wires (${m_{x}=m_{y}=N-c}$, $c=4$).
@@
In Fig. 4, we illustrate the scaling of the minimum gap for the different geometries of interest. Even though solving the MWIS problem for two vertices connected by an edge is trivial, within the gadget-encoding strategy, the adiabatic performance strongly depends on the considered geometry and target solution. Indeed, for the same graph problem, we observe both polynomially and exponentially closing minimum gaps, as shown in Fig. 4. For instance, in the (b) configuration, $\delta E_{m}$ closes polynomially if the target solution is $|01\rangle_{l}$ and exponentially if the target solution is $|10\rangle_{l}$: it is more challenging to reach the $|1\rangle_{l}$ state along the vertex-wire with the longest leg before the gadget.
@@ IV Perturbation theory and adiabatic performance with logical protocol
Figure 5: Schematic representation of the effective Hamiltonian. (a) (top) For a vertex-wire of $N$ atoms, the domain-wall states are characterized by two nearby atoms in the ground state (circled in gray), constituting the interface between two opposite $\mathbb{Z}_{2}$-ordered regions. They are labeled as $|j\rangle_{d}$, with $j=1,\dots,L=N/2+1$, according to the possible locations of the domain wall along the vertex-wire. (bottom) Illustration of the effective Hamiltonian in Eq. (4). Each domain-wall state is represented as a site of a 1D lattice. Links represent hopping terms between the connected sites; diagonal terms are represented via colors: white for zero, orange for $\epsilon_{i}$, and blue for $\epsilon_{f}$. (b,c) Illustration of the effective Hamiltonian in Eq. (5) for the (b) crossing and (c) crossing-with-edge setups. The two-domain-wall states, labeled by $|{\bm{j}}\rangle_{d}=|j_{x},j_{y}\rangle_{d}$, form a 2D lattice. Links represent hopping terms; diagonal terms are represented via colors: white for zero, orange for $\epsilon_{i,w}$, and blue for $\epsilon_{f,w}$ (with $w=x,y$). The sites at the corners of the lattice correspond to the logical states, and the diagonal terms on them are the sum of two boundary detuning (e.g., $\epsilon_{i,x}+\epsilon_{i,y}$ for $|00\rangle_{l}$). The insets show the states $|\zeta_{x},\zeta_{y}\rangle_{d}$ (for the crossing setup) and $|\zeta_{x},\zeta_{y}+1\rangle_{d}$ (for the crossing-with-edge setup).
In this section, we develop a perturbative description to gain insights into the dynamics and adiabatic performance of two intersecting vertex-wires, in the specific case of $\Delta_{g}=\Delta$. We develop an effective model within the lowest energy subspace of the Rydberg Hamiltonian in Eq. (2) for $\Omega,\epsilon\ll\Delta$, which consists of states with a single domain wall along each vertex-wire. This model allows us to understand the relationship between the geometry of the CWE setup and the minimum gap scaling along the logical protocol.
IV.1 Vertex-wire
We first discuss the case of a single vertex-wire [see Fig. 1(b)]. For $\Omega=\epsilon_{i}=\epsilon_{f}=0$, the lowest energy subspace of the Rydberg Hamiltonian in Eq. (2) is spanned by the so-called domain-wall states, shown in Fig. 5(a). The domain-wall states are defined by the presence of two neighboring atoms in the ground state, i.e., a domain wall, constituting the interface between two opposite $\mathbb{Z}_{2}$-ordered regions. We label these states as $|j\rangle_{d}$, with ${j=1,\dots,L=N/2+1}$, according to the possible domain-wall locations along the vertex-wire. We note that the two logical states correspond to $|1\rangle_{d}=|0\rangle_{l}$ and $|L\rangle_{d}=|1\rangle_{l}$.
@@ IV.2 Crossing setup
We now describe the crossing setup with an analogous formalism. The ground states of Eq. (2) for ${\Omega=\epsilon_{i,w}=\epsilon_{f,w}=0}$ are the product states between single-domain-wall states. They can thus be labeled as $|\bm{j}\rangle_{d}=|j_{x},j_{y}\rangle_{d}$, where ${j_{w}=1,\dots,L_{w}=N_{w}/2+2}$ corresponds to the position of the domain wall along the vertex-wire $w$ (with $w=x,y$). Thus, the ground states span a two-domain-wall Hilbert space, which can be represented as a two-dimensional (2D) rectangular lattice, whose corners correspond to the four logical states [see Fig. 5(b)].
@@ IV.2.1 Effective Hamiltonian
with hopping amplitude $t=\Omega^{2}/\Delta$. This equation is valid for $\epsilon_{i,w},\epsilon_{f,w}\ll\Delta$, and it is pictorially illustrated in Fig. 5(b). Here, $(\bm{j},\bm{j}^{\prime})$ denotes the sum over all pairs of states connected by a link, and $\epsilon(\bm{j})$ indicates a site-dependent diagonal term. Furthermore, we denote with $\zeta_{w}=m_{w}/2+1$ the last domain-wall position before the gadget along the vertex-wire $w$.
We note that the dynamics along the vertical and horizontal directions are not completely decoupled. Indeed, the gadget introduces two diagonal hopping matrix elements between $|\zeta_{x},\zeta_{y}\rangle_{d}$ and $|\zeta_{x}+1,\zeta_{y}+1\rangle_{d}$, and between $|\zeta_{x}+1,\zeta_{y}\rangle_{d}$ and $|\zeta_{x},\zeta_{y}+1\rangle_{d}$. These additional hoppings effectively lead to an interaction between domain walls inside the gadget. We further note that $\epsilon(\bm{j})$ is non-zero only on the boundaries of the domain-wall lattice (orange and blue circles), where its value can be controlled via local detuning on the extremes of the vertex-wires. In particular, the down, up, left, and right boundary lines are controlled via $\epsilon_{i,y}$, $\epsilon_{f,y}$, $\epsilon_{i,x}$, and $\epsilon_{f,x}$, respectively. At the four corners, the diagonal terms are the sum of those of the two adjacent boundary lines [e.g., $\epsilon(0,0)=\epsilon_{i,x}+\epsilon_{i,y}$].
@@ IV.2.2 Polynomial runtime via diabatic evolution
Figure 6: Analysis of the performance of the crossing setup, when considering the logical protocol. (a) Energy difference between the instantaneous eigenstates ($E_{n}$) and the instantaneous ground state ($E_{\rm GS}$) of the effective Hamiltonian, normalized to $t_{0}=\Omega_{0}^{2}/\Delta_{0}$ (where $\Omega_{0}$ and $\Delta_{0}$ are the maximum Rabi frequency and detuning), as a function of time $\tau$, when targeting $|11\rangle_{l}$ as the final ground state. We consider two vertex-wires of $N=28$ atoms intersecting at the middle ($m_{x}=m_{y}=N/2$). The spectrum is colored according to the overlap between the evolved state $|\psi(\tau)\rangle$ and the instantaneous eigenstates $|\phi_{n}(\tau)\rangle$ of the Hamiltonian: ${|\langle\phi_{n}(\tau)|\psi(\tau)\rangle|^{2}}$. The final time is set to $Tt_{0}=480$, such that $|\langle\psi(T)|11\rangle_{l}|^{2}\geq 0.9$. (b) Scaling of the algorithm runtime needed to achieve $|\langle\psi(T)|11\rangle_{l}|^{2}\geq 0.9$, as a function of $L=N/2+2$.
Since the crossing gadget has been designed to effectively decouple the logical degrees of freedom of two intersecting vertex-wires, we expect the crossing setup to behave as two independent vertex-wires during the adiabatic computation. This behavior is confirmed in Fig. 6(b), where we observe an $L^{2}$ scaling of the algorithm runtime for the logical protocol (see Fig. 2), when targeting $|11\rangle_{l}$ as the final ground state [with $\epsilon_{i,w}(0)=\epsilon_{f,w}(T)=2t_{0}$, where $w=x,y$ and $t_{0}=\Omega_{0}^{2}/\Delta_{0}$].
While this polynomial runtime is expected, it is nevertheless interesting to investigate the instantaneous energy spectrum of the system along the adiabatic protocol [see Fig. 6(a)]. Perhaps surprisingly, we observe two exponentially closing gaps. These gaps are a direct consequence of the exponentially small overlap between the ground state at half-protocol (i.e., at $\tau\simeq T/2$) and the initial and final ground states of the adiabatic protocol. Indeed, at half-protocol, the two-domain-wall ground-state wavefunction is localized around the gadget and decays exponentially along the legs of the vertex-wires, forming a so-called bound state. Despite these exponential gaps, the optimization problem is still solved in a polynomial runtime by exploiting a diabatic evolution. Indeed, at a polynomial timescale, the system does not couple to the unfavorable bound state; instead, it transitions from the instantaneous ground state to the first excited state at the first gap closing point and then returns to the ground state at the second one [71, 72, 5, 73, 74].
@@ IV.3 Crossing-with-edge setup
The crossing-with-edge gadget [see Fig. 1(d)] effectively imposes the $\mu_{x}\mu_{y}=0$ constraint between the logical states of two intersecting vertex-wires ($\mu_{w}=0,1$ with $w=x,y$). This constraint induces a mutual restriction on the position of the domain walls along the vertex-wires, allowing at most one domain wall to go through the gadget. As a result, the two-domain-wall Hilbert space acquires the $L$-shaped geometry depicted in Fig. 5(c), where, compared to the case of the crossing gadget, the lattice sites with both $j_{x}>\zeta_{x}$ and $j_{y}>\zeta_{y}$ are forbidden. This two-domain-wall space can be divided into three regions: one lead extending along $x$, $\left[\zeta_{x}+1,L_{x}\right]\times\left[0,\zeta_{y}\right]$; one lead extending along $y$, $\left[0,\zeta_{x}\right]\times\left[\zeta_{y}+1,L_{y}\right]$; and the inner region, $\left[0,\zeta_{x}\right]\times\left[0,\zeta_{y}\right]$. Here, $L_{w}=N_{w}/2+1$ with $w=x,y$.
In the $\Omega,\epsilon\ll\Delta$ limit, the dynamics are described by the effective Hamiltonian in Eq. (5), with links and diagonal terms represented in Fig. 5(c). This Hamiltonian is analogous to the one of the crossing setup, except for the removal of the sites in the forbidden region. Depending on whether or not the diagonal terms in Eq. (5) can be controlled by applying a detuning to the atoms at the extremes of the vertex-wires, we subdivide the boundary of the Hilbert space into a logical component [colored sites in Fig. 5(c)] and a non-logical one. In particular, the logical component consists of states where one (or both) vertex-wire is in a logical state.
@@ IV.3.1 Minimum gap scaling and domain-wall localization
Figure 7: Analysis of the performance of the crossing-with-edge setup for the four configurations in Fig. 4, when considering the logical protocol. (left) Ground-state wavefunction $|\mathrm{GS}(\bm{j})|^{2}=|\langle\bm{j}|\mathrm{GS}\rangle|^{2}$ at half-protocol ($\tau=T/2$), when ${\epsilon_{i,w}=\epsilon_{f,w}=0}$ for ${w={x,y}}$. For illustration purposes, we use $c=12$. (right) Scaling of the minimum gap $\delta E_{m}$, normalized to $t_{0}=\Omega_{0}^{2}/\Delta_{0}$ (where $\Omega_{0}$ and $\Delta_{0}$ are the maximum Rabi frequency and detuning) as a function of $L=N/2+1$ (with $N$ being the number of atoms per vertex-wire), setting $c=3$ (analogous to Fig. 4). For symmetric geometries (a,c,d), the ground state at final time is $|01\rangle_{l}$. In the asymmetric geometry (b), $|10\rangle_{l}$ (blue) and $|01\rangle_{l}$ (orange) are considered separately.
As observed in Sec. III, in the CWE setup, the minimum gap scaling strongly depends on the geometry of the system and target state. Here, we aim to clarify this dependence by studying the two-domain-wall dynamics along the logical protocol, for the same geometrical configurations considered in the standard protocol (see Fig. 4). In Fig. 7, we show (right) the scaling of the minimum gap and (left) the ground-state wavefunction at half-protocol, where ${\epsilon_{i,w}=\epsilon_{f,w}=0}$. Specifically, we consider two vertex-wires with the same number of atoms, i.e., $L_{x}=L_{y}=L=N/2+1$, and perform the logical protocol of Fig. 2 by choosing: ${\epsilon_{i,x}(0)=\epsilon_{i,y}(0)=2t_{0}}$, ${\epsilon_{f,x}(0)=\epsilon_{f,y}(0)=0}$, ${\epsilon_{i,x}(T)=\epsilon_{i,y}(T)=4t_{0}}$ and $\epsilon_{f,x}(T)=\frac{12}{5}\epsilon_{f,y}(T)=12t_{0}$ ($\epsilon_{f,y}(T)=\frac{12}{5}\epsilon_{f,x}(T)=12t_{0}$) to target $|10\rangle_{l}$ ($|01\rangle_{l}$) as final ground state.
As for the standard protocol, we observe both polynomial and exponential scaling, depending on the geometry and the target state. In particular, $\delta E_{m}$ scales polynomially in configurations (a, d), exponentially in (c), and in (b), depending on the target state, it exhibits both behaviors. We can predict whether the scaling is exponential or polynomial based on the overlap between the ground-state wavefunction before the minimum gap point and the final target state [75, 76, 77, 78, 79]. In particular, when the wavefunction at half-protocol decays exponentially in the lead where the target state is located — as in (c) and in (b) for the target state $|10\rangle_{l}$ — we find an exponential scaling of the minimum gap. Here, in contrast to the crossing setup, there is only a single exponentially closing gap dominating the dynamics. Therefore, diabatic transitions to higher excited states cannot be exploited to speed up the algorithm runtime.
@@
with $\bm{r}=(x,y)$ being defined in a continuous $L$-shaped domain, and $k^{2}\simeq E_{k}/t+4$ for $k\ll 1$. The boundary conditions of Eq. (6) are determined by $\epsilon({\bm{j}})$. At ${\tau=T/2}$, $\epsilon({\bm{j}})=0$ maps to homogeneous Dirichlet boundary conditions (DBC). In the following, we use this model to infer the ground-state wavefunction in the (b) and (c) configurations.
@@
In the (c) configuration, populating the leads would require a high kinetic cost, scaling as ${\sim(\pi/c)^{2}}$. The domain-wall wavefunction thus minimizes the kinetic energy by localizing in the inner region. This effectively traps the domain walls before the gadget, forming a so-called bound state. The emergence of such a bound state in an $L$-shaped Hilbert space is a purely quantum phenomenon, which was first noted for quark confinement and hadronic interactions [80], and extensively studied in the field of quantum waveguides [81, 82, 83]. This bound state is separated from the continuum spectrum by a finite energy gap $\delta E$, which does not close in the infinite-$L$ limit. In Appendix E, we analytically describe such phenomenon within the so-called single-mode approximation and via conformal mapping methods. We predict the energy gap and compute the geometric range of $\zeta_{x}/\zeta_{y}$ for which it remains finite in the $L\to\infty$ limit.
@@
Figure 8: Bound state analysis in the non-perturbative regime. Parameters of the gap scaling in Eq. (7) for ${\zeta=3}$ as a function of $\Omega/\Delta$. The dashed gray line indicates the transition point known for one-dimensional Rydberg chains with nearest-neighbor blockade, $\Omega/\Delta\sim 1/1.3$ [49].

### v2

@@ III.2 Crossing-with-edge setup
Figure 4: Scaling of the minimum gap $\delta E_{m}$ along the standard protocol for the crossing-with-edge setup. $\delta E_{m}$ (normalized to the maximum Rabi frequency $\Omega_{0}$) as a function of the number of atoms per vertex wire $N$, for the configurations: (a) $m_{x}=m_{y}=N/2$, (b) $m_{x}=N/2$, $m_{y}=4$, (c) $m_{x}=m_{y}=4$, and (d) $m_{x}=m_{y}=N-4$. For symmetric geometries (a,c,d), the ground state at final time is $|01\rangle_{l}$. In the asymmetric geometry (b), $|10\rangle_{l}$ (blue) and $|01\rangle_{l}$ (orange) are considered separately. In the top-right inset of (c), the gap $\delta E$ is shown as a function of $\Delta/\Omega$, for $N$ ranging from $8$ (light blue) to $48$ (dark blue).
Two vertex wires intersecting via the CWE gadget [see Fig. 1(d)] encode the simple graph consisting of two vertices connected by an edge. Even though the two vertices play a symmetric role in the original graph, this is not necessarily the case for the two vertex wires if $m_{x}\neq m_{y}$. We thus consider, for a given geometry (i.e., for a given choice of $m_{x}$ and $m_{y}$), two distinct MWIS problems, where $\delta_{x}>\delta_{y}$ and $\delta_{x}<\delta_{y}$. The solution to these problems is encoded in the $|10\rangle_{l}$ and $|01\rangle_{l}$ states of the two vertex wires, respectively. Motivated by the possible positions of the gadget in a large network, we analyze four different geometrical configurations with $N_{x}=N_{y}=N$. We consider the gadget placed: (a) at the midpoint of the two vertex wires (symmetric case, ${m_{x}=m_{y}=N/2}$); (b) close to the beginning of the $y$ vertex wire ($m_{x}=N/2$, $m_{y}=c=4$); (c) close to the beginning of both vertex wires ($m_{x}=m_{y}=c=4$); (d) close to the end of both vertex wires (${m_{x}=m_{y}=N-c}$, $c=4$).
@@
In Fig. 4, we illustrate the scaling of the minimum gap for the different geometries of interest. Even though solving the MWIS problem for two vertices connected by an edge is trivial, within the gadget-encoding strategy, the adiabatic performance strongly depends on the considered geometry and target solution. Indeed, for the same graph problem, we observe both polynomially and exponentially closing minimum gaps, as shown in Fig. 4. For instance, in the (b) configuration, $\delta E_{m}$ closes polynomially if the target solution is $|01\rangle_{l}$ and exponentially if the target solution is $|10\rangle_{l}$: it is more challenging to reach the $|1\rangle_{l}$ state along the vertex wire with the longest leg before the gadget.
@@ IV Perturbation theory and adiabatic performance with logical protocol
Figure 5: Schematic representation of the effective Hamiltonian. (a) (Top) For a vertex wire of $N$ atoms, the domain-wall states are characterized by two nearby atoms in the ground state (circled in gray), constituting the interface between two opposite $\mathbb{Z}_{2}$-ordered regions. They are labeled as $|j\rangle_{d}$, with $j=1,\dots,L=N/2+1$, according to the possible locations of the domain wall along the vertex wire. (Bottom) Illustration of the effective Hamiltonian in Eq. (4). Each domain-wall state is represented as a site of a 1D lattice. Links represent hopping terms between the connected sites; diagonal terms are represented via colors: white for zero, orange for $\epsilon_{i}$, and blue for $\epsilon_{f}$. (b,c) Illustration of the effective Hamiltonian in Eq. (5) for the (b) crossing and (c) crossing-with-edge setups. The two-domain-wall states, labeled by $|{\bm{j}}\rangle_{d}=|j_{x},j_{y}\rangle_{d}$, form a 2D lattice. Links represent hopping terms; diagonal terms are represented via colors: white for zero, orange for $\epsilon_{i,w}$, and blue for $\epsilon_{f,w}$ (with $w=x,y$). The sites at the corners of the lattice correspond to the logical states, and the diagonal terms on them are the sum of two boundary detuning (e.g., $\epsilon_{i,x}+\epsilon_{i,y}$ for $|00\rangle_{l}$). The insets show the states $|\zeta_{x},\zeta_{y}\rangle_{d}$ (for the crossing setup) and $|\zeta_{x},\zeta_{y}+1\rangle_{d}$ (for the crossing-with-edge setup).
In this section, we develop a perturbative description to gain insights into the dynamics and adiabatic performance of two intersecting vertex wires, in the specific case of $\Delta_{g}=\Delta$. We develop an effective model within the lowest-energy subspace of the Rydberg Hamiltonian in Eq. (2) for $\Omega,\epsilon\ll\Delta$, which consists of states with a single domain wall along each vertex wire. This model allows us to understand the relationship between the geometry of the CWE setup and the minimum gap scaling along the logical protocol.
IV.1 Vertex wire
We first discuss the case of a single vertex wire [see Fig. 1(b)]. For $\Omega=\epsilon_{i}=\epsilon_{f}=0$, the lowest-energy subspace of the Rydberg Hamiltonian in Eq. (2) is spanned by the so-called domain-wall states, shown in Fig. 5(a). The domain-wall states are defined by the presence of two neighboring atoms in the ground state, i.e., a domain wall, constituting the interface between two opposite $\mathbb{Z}_{2}$-ordered regions. We label these states as $|j\rangle_{d}$, with ${j=1,\dots,L=N/2+1}$, according to the possible domain-wall locations along the vertex wire. We note that the two logical states correspond to $|1\rangle_{d}=|0\rangle_{l}$ and $|L\rangle_{d}=|1\rangle_{l}$.
@@ IV.2 Crossing setup
We now describe the crossing setup with an analogous formalism. The ground states of Eq. (2) for ${\Omega=\epsilon_{i,w}=\epsilon_{f,w}=0}$ are the product states between single-domain-wall states. They can thus be labeled as $|\bm{j}\rangle_{d}=|j_{x},j_{y}\rangle_{d}$, where ${j_{w}=1,\dots,L_{w}=N_{w}/2+2}$ corresponds to the position of the domain wall along the vertex wire $w$ (with $w=x,y$). Thus, the ground states span a two-domain-wall Hilbert space, which can be represented as a two-dimensional (2D) rectangular lattice, whose corners correspond to the four logical states [see Fig. 5(b)].
@@ IV.2.1 Effective Hamiltonian
with hopping amplitude $t=\Omega^{2}/\Delta$. This equation is valid for $\epsilon_{i,w},\epsilon_{f,w}\ll\Delta$, and it is pictorially illustrated in Fig. 5(b). Here, $(\bm{j},\bm{j}^{\prime})$ denotes the sum over all pairs of states connected by a link, and $\epsilon(\bm{j})$ indicates a site-dependent diagonal term. Furthermore, we denote with $\zeta_{w}=m_{w}/2+1$ the last domain-wall position before the gadget along the vertex wire $w$.
We note that the dynamics along the vertical and horizontal directions are not completely decoupled. Indeed, the gadget introduces two diagonal hopping matrix elements between $|\zeta_{x},\zeta_{y}\rangle_{d}$ and $|\zeta_{x}+1,\zeta_{y}+1\rangle_{d}$, and between $|\zeta_{x}+1,\zeta_{y}\rangle_{d}$ and $|\zeta_{x},\zeta_{y}+1\rangle_{d}$. These additional hoppings effectively lead to an interaction between domain walls inside the gadget. We further note that $\epsilon(\bm{j})$ is nonzero only on the boundaries of the domain-wall lattice (orange and blue circles), where its value can be controlled via local detuning on the extremes of the vertex wires. In particular, the down, up, left, and right boundary lines are controlled via $\epsilon_{i,y}$, $\epsilon_{f,y}$, $\epsilon_{i,x}$, and $\epsilon_{f,x}$, respectively. At the four corners, the diagonal terms are the sum of those of the two adjacent boundary lines [e.g., $\epsilon(0,0)=\epsilon_{i,x}+\epsilon_{i,y}$].
@@ IV.2.2 Polynomial runtime via diabatic evolution
Figure 6: Analysis of the performance of the crossing setup, when considering the logical protocol. (a) Energy difference between the instantaneous eigenstates ($E_{n}$) and the instantaneous ground state ($E_{\rm GS}$) of the effective Hamiltonian, normalized to $t_{0}=\Omega_{0}^{2}/\Delta_{0}$ (where $\Omega_{0}$ and $\Delta_{0}$ are the maximum Rabi frequency and detuning), as a function of time $\tau$, when targeting $|11\rangle_{l}$ as the final ground state. We consider two vertex wires of $N=28$ atoms intersecting at the middle ($m_{x}=m_{y}=N/2$). The spectrum is colored according to the overlap between the evolved state $|\psi(\tau)\rangle$ and the instantaneous eigenstates $|\phi_{n}(\tau)\rangle$ of the Hamiltonian: ${|\langle\phi_{n}(\tau)|\psi(\tau)\rangle|^{2}}$. The final time is set to $Tt_{0}=480$, such that $|\langle\psi(T)|11\rangle_{l}|^{2}\geq 0.9$. (b) Scaling of the algorithm runtime needed to achieve $|\langle\psi(T)|11\rangle_{l}|^{2}\geq 0.9$, as a function of $L=N/2+2$.
Since the crossing gadget has been designed to effectively decouple the logical degrees of freedom of two intersecting vertex wires, we expect the crossing setup to behave as two independent vertex wires during the adiabatic computation. This behavior is confirmed in Fig. 6(b), where we observe an $L^{2}$ scaling of the algorithm runtime for the logical protocol (see Fig. 2), when targeting $|11\rangle_{l}$ as the final ground state [with $\epsilon_{i,w}(0)=\epsilon_{f,w}(T)=2t_{0}$, where $w=x,y$ and $t_{0}=\Omega_{0}^{2}/\Delta_{0}$].
While this polynomial runtime is expected, it is nevertheless interesting to investigate the instantaneous energy spectrum of the system along the adiabatic protocol [see Fig. 6(a)]. Perhaps surprisingly, we observe two exponentially closing gaps. These gaps are a direct consequence of the exponentially small overlap between the ground state at half-protocol (i.e., at $\tau\simeq T/2$) and the initial and final ground states of the adiabatic protocol. Indeed, at half-protocol, the two-domain-wall ground-state wavefunction is localized around the gadget and decays exponentially along the legs of the vertex wires, forming a so-called bound state. Despite these exponential gaps, the optimization problem is still solved in a polynomial runtime by exploiting a diabatic evolution. Indeed, at a polynomial timescale, the system does not couple to the unfavorable bound state; instead, it transitions from the instantaneous ground state to the first excited state at the first gap closing point and then returns to the ground state at the second one [73, 74, 5, 75, 76].
@@ IV.3 Crossing-with-edge setup
The crossing-with-edge gadget [see Fig. 1(d)] effectively imposes the $\mu_{x}\mu_{y}=0$ constraint between the logical states of two intersecting vertex wires ($\mu_{w}=0,1$ with $w=x,y$). This constraint induces a mutual restriction on the position of the domain walls along the vertex wires, allowing at most one domain wall to go through the gadget. As a result, the two-domain-wall Hilbert space acquires the $L$-shaped geometry depicted in Fig. 5(c), where, compared to the case of the crossing gadget, the lattice sites with both $j_{x}>\zeta_{x}$ and $j_{y}>\zeta_{y}$ are forbidden. This two-domain-wall space can be divided into three regions: one lead extending along $x$, $\left[\zeta_{x}+1,L_{x}\right]\times\left[0,\zeta_{y}\right]$; one lead extending along $y$, $\left[0,\zeta_{x}\right]\times\left[\zeta_{y}+1,L_{y}\right]$; and the inner region, $\left[0,\zeta_{x}\right]\times\left[0,\zeta_{y}\right]$. Here, $L_{w}=N_{w}/2+1$ with $w=x,y$.
In the $\Omega,\epsilon\ll\Delta$ limit, the dynamics are described by the effective Hamiltonian in Eq. (5), with links and diagonal terms represented in Fig. 5(c). This Hamiltonian is analogous to the one of the crossing setup, except for the removal of the sites in the forbidden region. Depending on whether or not the diagonal terms in Eq. (5) can be controlled by applying a detuning to the atoms at the extremes of the vertex wires, we subdivide the boundary of the Hilbert space into a logical component [colored sites in Fig. 5(c)] and a nonlogical one. In particular, the logical component consists of states where one (or both) vertex wire is in a logical state.
@@ IV.3.1 Minimum gap scaling and domain-wall localization
Figure 7: Analysis of the performance of the crossing-with-edge setup for the four configurations in Fig. 4, when considering the logical protocol. (Left) Ground-state wavefunction $|\mathrm{GS}(\bm{j})|^{2}=|\langle\bm{j}|\mathrm{GS}\rangle|^{2}$ at half-protocol ($\tau=T/2$), when ${\epsilon_{i,w}=\epsilon_{f,w}=0}$ for ${w={x,y}}$. For illustration purposes, we use $c=12$. (Right) Scaling of the minimum gap $\delta E_{m}$, normalized to $t_{0}=\Omega_{0}^{2}/\Delta_{0}$ (where $\Omega_{0}$ and $\Delta_{0}$ are the maximum Rabi frequency and detuning) as a function of $L=N/2+1$ (with $N$ being the number of atoms per vertex wire), setting $c=3$ (analogous to Fig. 4). For symmetric geometries (a,c,d), the ground state at final time is $|01\rangle_{l}$. In the asymmetric geometry (b), $|10\rangle_{l}$ (blue) and $|01\rangle_{l}$ (orange) are considered separately.
As observed in Sec. III, in the CWE setup, the minimum gap scaling strongly depends on the geometry of the system and target state. Here, we aim to clarify this dependence by studying the two-domain-wall dynamics along the logical protocol, for the same geometrical configurations considered in the standard protocol (see Fig. 4). In Fig. 7, we show (right) the scaling of the minimum gap and (left) the ground-state wavefunction at half-protocol, where ${\epsilon_{i,w}=\epsilon_{f,w}=0}$. Specifically, we consider two vertex wires with the same number of atoms, i.e., $L_{x}=L_{y}=L=N/2+1$, and perform the logical protocol of Fig. 2 by choosing: ${\epsilon_{i,x}(0)=\epsilon_{i,y}(0)=2t_{0}}$, ${\epsilon_{f,x}(0)=\epsilon_{f,y}(0)=0}$, ${\epsilon_{i,x}(T)=\epsilon_{i,y}(T)=4t_{0}}$ and $\epsilon_{f,x}(T)=\frac{12}{5}\epsilon_{f,y}(T)=12t_{0}$ ($\epsilon_{f,y}(T)=\frac{12}{5}\epsilon_{f,x}(T)=12t_{0}$) to target $|10\rangle_{l}$ ($|01\rangle_{l}$) as the final ground state.
As for the standard protocol, we observe both polynomial and exponential scaling, depending on the geometry and the target state. In particular, $\delta E_{m}$ scales polynomially in configurations (a), (d), exponentially in (c), and in (b), depending on the target state, it exhibits both behaviors. We can predict whether the scaling is exponential or polynomial based on the overlap between the ground-state wavefunction before the minimum gap point and the final target state [77, 78, 79, 80, 81]. In particular, when the wavefunction at half-protocol decays exponentially in the lead where the target state is located — as in (c) and in (b) for the target state $|10\rangle_{l}$ — we find an exponential scaling of the minimum gap. Here, in contrast to the crossing setup, there is only a single exponentially closing gap dominating the dynamics. Therefore, diabatic transitions to higher excited states cannot be exploited to speed up the algorithm runtime.
@@
with $\bm{r}=(x,y)$ being defined in a continuous $L$-shaped domain, and $k^{2}\simeq E_{k}/t+4$ for $k\ll 1$. The boundary conditions of Eq. (6) are determined by $\epsilon({\bm{j}})$. At ${\tau=T/2}$, $\epsilon({\bm{j}})=0$ maps to homogeneous Dirichlet boundary conditions (DBCs). In the following, we use this model to infer the ground-state wavefunction in the (b) and (c) configurations.
@@
In the (c) configuration, populating the leads would require a high kinetic cost, scaling as ${(\pi/c)^{2}}$. The domain-wall wavefunction thus minimizes the kinetic energy by localizing in the inner region. This effectively traps the domain walls before the gadget, forming a so-called bound state. The emergence of such a bound state in an $L$-shaped Hilbert space is a purely quantum phenomenon, which was first noted for quark confinement and hadronic interactions [82], and extensively studied in the field of quantum waveguides [83, 84, 85]. This bound state is separated from the continuum spectrum by a finite energy gap $\delta E$, which does not close in the infinite-$L$ limit. In Appendix E, we analytically describe such phenomenon within the so-called single-mode approximation and via conformal mapping methods. We predict the energy gap and compute the geometric range of $\zeta_{x}/\zeta_{y}$ for which it remains finite in the $L\to\infty$ limit.
@@
Figure 8: Bound-state analysis in the nonperturbative regime. Parameters of the gap scaling in Eq. (7) for ${\zeta=3}$ as a function of $\Omega/\Delta$. The dashed gray line indicates the transition point known for one-dimensional Rydberg chains with nearest-neighbor blockade, $\Omega/\Delta\sim 1/1.3$ [51].